Today, Bruce and Brenda each have $150,000 in an investment account. No other contributions will be made to their investment accounts. Both have the same goal: They each want their account to reach S1 million, at which time each will retire. Bruce has his money invested in risk-free securities with an expected monthly return of 0.41 percent. Brenda has her money invested in a stock fund with an expected semi-annual return of 4.9 percent. How many years after Brenda retires will Bruce retire? Choose the closest response. a. 12.6 о 19.0 c. 19.9 d. 29.4 e. Insufficient Information

How to find EAR1 and EAR2?   

How to find the years ?
Answer #1

SEE THE SCREENSHOT. ANY DOUBTS, HAPPY TO HELP YOU. THANK YOU. THUMBS UP PLEASE.

TO MAKE EAR MORE CLEAR, I HAVE USED FORMULA FOR CALCULATING EAR. FOR SECOND PART, USED EXCEL. THANK YOU

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text aCopy B l u. ,_a. ars-函Merge & Center, $, % , 弼,8 C Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard Alignment Number Cells CT194 CS CT CU CV CW CX CY CZ DA DB DC PART 1 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 /4 | CALCULATOR . LOAN OPTIONS EAR = (1 +r/m)^m-1 MONTHLY RETURN 0.41% so r/m-0.41% AND m-12 EAR = (1 +0.41%)^12-1- 5.0325% SEMI ANNUAL RETURN 4.90% so r/m = 4.9% AND m-2 EARs (1 +4.996)^2-1- 10.0401% CAP STRU VALUE BOX DILUTION DOLLAR COST AVGMORTGAGE EXPO Sheeti Sheet2 ShelI rences: x261 福 130% 19-01-2019Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy ▼ в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard Alignment Number Cells Edting CY197 CS CT CU CV CX CY CZ DA DB DC 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 /4 | CALCULATOR . LOAN OPTIONS NOW TO FIND N, WE WILL USE EXCEL FUNCTION NPER BRUCE PV FV RATE NO OF YEARS 150000 1000000 5.0325% 38.64 NPER(CU198,-CU196,CU197) BRENDA PV FV RATE NO OF YEARS 150000 1000000 10.0401% 19.83 NPER(CU204,,-CU202,CU203) So BRUCE WILL TAKE 19 YEARS MORE TO RETIRE (38.64 - 19.83) AS BRUCE RATE OF RETURN IS LOWER, HE WILL TAKE MORE YEARS TO REACH 1 MILLION AND HIS RATE IS AROUND HALF THAN THAT OF BRENDA SO WILL TAKE DOUBLE TIME THAN BRENDA TO MAKE MILLION CAP STRU VALUE BOX DILUTION DOLLAR COST AVGMORTGAGE EXPO Sheeti Sheet2 ShelI rences: x261 福 130% 05:13 19-01-2019